OpenStudy (anonymous):

*Medal* 1) −4x − 2y = −12 4x + 8y = −24

OpenStudy (anonymous):

I have to solve through Elimination

OpenStudy (anonymous):

do you HAVE to or can you use substitution?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

in columns

OpenStudy (anonymous):

\[ −4x − 2y = −12 \\ 4x + 8y = −24\]_________________

OpenStudy (anonymous):

what do you get on the right when you add?

OpenStudy (anonymous):

in other words, what is \(-12+(-24)\) ?

OpenStudy (austinl):

-4x+4x=? -2y+8y=? -12+(-24)=?

OpenStudy (anonymous):

-12-24 = -36

OpenStudy (anonymous):

ok so the right hand side is \(-36\) when you add how about the left hand side? what is \(-4x+4x\) ?

OpenStudy (anonymous):

Nothing, Zero, am I right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the \(x\) terms add up to zero what is\(-2y+8y\) ?

OpenStudy (anonymous):

+4y

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

\(8-2=?\)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

ok so \(8y-2y=6y\) right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

therefore if you add you get the new equation \[6y=-36\] one equation only and also only one variable if \(6y=-36\) then what is \(y\) ?

OpenStudy (anonymous):

y = -6 right?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Then you should be able to plug that in for y in one of the original equations and solve for x. Or alternatively, you could go through the elimination process again...